OpenStudy (anonymous):

The price of a storybook is RM2 more than that of a comic book.Encik Ahamd buys x storybooks at a price of RM108.With the same amount of money,he can buy (x+9) comic books. a)Form a quadratic equation for the price of the storybooks and comic books in terms of x and show that it can be simplified to x^2+9x-486=0

OpenStudy (anonymous):

@ganeshie8

OpenStudy (michele_laino):

sorry if I say that price of storybook is RM108 and price of comicbook is RM106, is it right?

OpenStudy (michele_laino):

please @ganeshie8

OpenStudy (michele_laino):

or please @MARC_

OpenStudy (michele_laino):

if it is so, then we can write: 108x=106x+9*106 then x=477

OpenStudy (michele_laino):

sorry, I've made an error!

OpenStudy (michele_laino):

I got! if we factorize your quadratic equation, we will get this: \[x ^{2}+9x-486=(x-18)(x+27)=0\] so the unique acceptable solution is x=18. now from the textof your problem if I call with s the price of storybook, and c the price of comicbook, we can write; s=c+2 then: (c+2)*x=c*(x+9) so: 2*x=9*c finally inserting x=18, in the last above equation, we will get: c=4, and s=6

OpenStudy (michele_laino):

oops c=RM4 and s=RM6

OpenStudy (michele_laino):

@MARC_